OpenStudy (anonymous):

Proof by induction. I would like to please have help on how to prove that: 2^n < n! for n > 4

OpenStudy (anonymous):

it is true for n = 5 since \[2^5=32\] whereas \[5!=120\] now assume true for all \[k\leq n\] prove it is true for n + 1

OpenStudy (anonymous):

that is prove \[2^{n+1}<(n+1)!\] i.e. \[2\times 2^n<(n+1)n!\] and since BY INDUCTION \[2^n<n!\] and \[2<n+1\] we know \[2\times 2^n<(n+1)n!\]i.e. \[2^{n+1}<(n+1)!\]

OpenStudy (anonymous):

makes perfect sense :) thank you